OpenStudy (anonymous):

More SAT practice. Very confused. What is the average (arithmetic mean) of all the multiples of ten from 10 to 190 inclusive? A. 90 B. 95 C. 100 D. 105 E. 110 Explain to me what in inclusive means?

OpenStudy (igreen):

An inclusive range of numbers includes the first and last number and all numbers in between..

OpenStudy (anonymous):

o so 10 through 190

OpenStudy (igreen):

First, we calculate the number of terms.. 10, 20, 30, 40, 50, 60, 70, 80, 90, 100, 110, 120, 130, 140, 150, 160, 170, 180, 190. That's 19.

OpenStudy (igreen):

Plug it in our sequence formula: \(S = (\dfrac{19}{2})(a + l)\) Plug in what we know: \(S = (\dfrac{19}{2})(10 + 190)\) Can you solve that?

OpenStudy (anonymous):

S= 1900?

OpenStudy (igreen):

Yep, now we divide that by the number of terms..which is 19. Can you do that?

OpenStudy (anonymous):

C. 100

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

Thank you for your help again